OpenStudy (drummerboyrevived):

37 - 12 = c - 5 c = __

OpenStudy (anonymous):

\[37-12=c-5\]\[25=c-5\]\[30=c\] To check:\[37-12=(30)-5\]\[25=25\] CHECK